# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v3

@@ Conclusion.
Combining all bounds:
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$
For SHAKE-256 as PRF, $\epsilon_{\mathsf{PRF}}$ is negligible, yielding the stated bound.
@@ 4.3 Privacy
Theorem 4.4 (Privacy).
If PRF is secure, then the view of any coalition $C$ with $|C|<T$ can be efficiently simulated given only $\mathsf{pk}$, corrupted shares, and final signatures.
@@
We construct simulator $\mathcal{S}$ and prove indistinguishability via hybrids.
Simulator $\mathcal{S}$.
Given $\mathsf{pk}$, corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$, and signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$:
  1. 1.
    Round 1: For each honest $j$, sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$.
  2. 2.
    Round 2: Compute $\mathbf{w}_{1}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$. For each honest $j$, sample $\mathbf{W}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ subject to $\sum_{j\notin C}\mathbf{W}_{j}=\mathbf{w}-\sum_{i\in C}\mathbf{W}_{i}$.
  3. 3.
    Round 3: Sample honest $\mathbf{U}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ subject to $\sum_{j\notin C}\mathbf{U}_{j}=\mathbf{z}-\sum_{i\in C}\mathbf{U}_{i}$.
Hybrid $\mathsf{H}_{0}$ (Real).
Real protocol execution with honest parties using true secrets.
Hybrid $\mathsf{H}_{1}$.
Replace PRF outputs for honest-honest pairs with truly random: for $j,k\notin C$, set $\mathsf{PRF}(\mathsf{seed}_{j,k},\cdot)\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{d}$. By PRF security: $|\Pr[\mathsf{H}_{1}]-\Pr[\mathsf{H}_{0}]|\leq\binom{|H|}{2}\cdot\epsilon_{\mathsf{PRF}}$ where $H=S\setminus C$.
Hybrid $\mathsf{H}_{2}$ (Simulated).
Replace honest contributions with $\mathcal{S}$’s output. Since $|H|\geq 2$, honest parties’ masks include at least one random term unknown to $\mathcal{A}$. By the one-time pad property over $R_{q}$: $\Pr[\mathsf{H}_{2}]=\Pr[\mathsf{H}_{1}]$.
Conclusion.
$\mathcal{S}$’s output is computationally indistinguishable from the real view with advantage $\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
Lemma 4.6 (Mask Hiding).
@@ Conclusion.
An honest party’s masked response $\mathbf{U}_{j}=\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}$ is computationally indistinguishable from uniform in $R_{q}^{\ell}$, given all information available to a coalition of <T<T parties.
Proof 4.7.
@@
Since $|S|\geq T+1$ and $|C|\leq T-1$, at least two honest parties exist. Party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{nonce})$ for some honest $k^{\prime}\neq j$. This seed is unknown to $\mathcal{A}$, so by PRF security, the term is indistinguishable from uniform in $R_{q}^{\ell}$. Thus $\mathbf{U}_{j}=(\text{fixed})+(\text{uniform})\equiv\text{uniform}$.
@@ 4.4 Nonce Distribution Security
Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$ where $R_{2}^{\epsilon}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers. Concretely for ML-DSA-65: $R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, and $\epsilon<10^{-30}$.
The square root arises from the generic Rényi divergence probability bound; specialized reductions for Fiat-Shamir with Aborts [Raccoon2024] can achieve tighter (linear) bounds. With ML-DSA-65’s 192-bit security ($\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$), the bound gives $\mathsf{Adv}_{\Pi}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$, maintaining adequate security margin. Detailed calculations appear in Appendix D.
@@ 4.6 Concrete Security Bounds
Share Extraction M-LWE + PRF 192 bits
Table 3: Concrete security of masked threshold ML-DSA-65. Parameters match NIST Security Level 3.
The threshold construction inherits the security level of single-signer ML-DSA with no meaningful additional loss.
@@ 4.7.1 Profile P1 UC Security
  • Input: Receive masked shares $\{(\mathbf{U}_{i},\mathbf{V}_{i})\}_{i\in S}$ from coordinator
@@
      Aggregate $\mathbf{z}=\sum_{i}\mathbf{U}_{i}\bmod q$ and $c\mathbf{s}_{2}=\sum_{i}\mathbf{V}_{i}\bmod q$
@@
  • Output: Send $(\tilde{c},\mathbf{z},\mathbf{h})$ or $\bot$ to all parties; never reveal $c\mathbf{s}_{2}$
@@
    $\mathcal{S}$ simulates honest parties’ Round 1–3 messages using the privacy simulator from Theorem 4.4 (masked values appear uniform).
@@
    $\mathcal{S}$ forwards all messages to/from $\mathcal{F}_{\mathsf{TEE}}$.
@@
    $\mathcal{F}_{\mathsf{TEE}}$ computes internally and outputs only the final signature (or $\bot$).
Since $\mathcal{F}_{\mathsf{TEE}}$ never reveals $c\mathbf{s}_{2}$ and the masked shares are indistinguishable from uniform (Lemma 4.6), the real and ideal executions are indistinguishable. The distinguishing advantage is bounded by $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (from the privacy proof).
@@ 4.7.2 Profile P3+ Semi-Async Security Model
Theorem 4.14 (Profile P3+ UC Security).
@@ 4.7.3 Summary of UC Security
  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.10). EUF-CMA and privacy follow from Theorems 4.2 and $4.$ Achieves 3 rounds.
@@ 5.2 Benchmark Methodology
Our Rust implementation ($\approx$4000 lines) provides optimized performance through parallel polynomial operations using Rayon, NTT-based polynomial multiplication with Montgomery reduction, and efficient EdaBit and Beaver triple pooling with lazy generation.
@@
All measurements are per-attempt: a single execution of the 3-round signing protocol, which either succeeds (outputs $\sigma$) or aborts (due to rejection sampling).

### v4

@@ Conclusion.
Games 4 and the M-SIS reduction give $\Pr[\mathsf{Win}_{3.5}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$. Composing with the Rényi hop (Game $3\to 3.5$) and the birthday bound (Game $1\to 2$):
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
For ML-DSA-65 with $|S|\leq 17$: $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$, $n\ell=1280$), $\epsilon_{\mathsf{IH}}<10^{-30}$, and SHAKE-256 yields negligible $\epsilon_{\mathsf{PRF}}$, giving the stated bound.
@@ 4.3 Privacy
Theorem 4.4 (Statistical Share Privacy).
With Shamir nonce DKG, for any signing set $S$ with $|S|\geq T$ and any coalition $C\subset S$ with $|C|\leq T-1$: each honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy
$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
where $\mathsf{View}_{C}$ denotes the adversary’s view: corrupted shares, DKG evaluations, challenge $c$, the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$, and all signatures (current and prior). Note that $\mathsf{View}_{C}$ does not include individual $\mathbf{W}_{i}$ values, which are sent only to the coordinator (Remark 3.7). Since $\mathbf{w}_{1}$ is already implicit in any valid signature, it adds no information beyond $\sigma$ itself. For ML-DSA-65 ($n\ell=1280$, $\gamma_{1}=2^{19}$), the bound exceeds $1280\cdot(20-\log_{2}|S|)$ bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 1280\cdot\log_{2}9\approx 4058$ bits for $|S|\leq 17$. No computational assumptions are needed; the privacy is statistical.
@@
The adversary’s knowledge about the honest party $h$’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the nonce DKG. Since $|C|\leq T-1$ and $\mathbf{f}_{h}$ has degree T1T-1 (hence $T$ coefficients), the adversary has at most T1T-1 linear equations in $T$ unknowns. The solution space has dimension $T-|C|\geq 1$ (exactly $1$ in the worst case |C|=T1|C|=T-1).
Null space argument. The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$. The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
Posterior constraint on $\mathbf{t}$. The coefficient $\mathbf{a}_{h,T-1}$ is sampled $R_{q}^{\ell}$-uniformly a priori. However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy the bounded-nonce constraint $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$. Since $\mathbf{n}(0)=\prod_{j\in C}(-x_{j})\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values. The posterior distribution of $\mathbf{t}$ is therefore not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of size $2\lfloor\gamma_{1}/|S|\rfloor+1$.
Min-entropy bound. Since $\mathbf{n}(x_{h})=\prod_{j\in C}(x_{h}-x_{j})\neq 0\pmod{q}$ (because $q$ is prime and all evaluation points are distinct), multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$. The honest party’s evaluation $\mathbf{f}_{h}(x_{h})=\mathbf{g}(x_{h})+\mathbf{t}\cdot\mathbf{n}(x_{h})$ therefore inherits the same per-coefficient support size. The nonce share $\mathbf{y}_{h}=\sum_{j\in C}\mathbf{f}_{j}(x_{h})+\mathbf{f}_{h}(x_{h})$ shifts by a known constant, preserving support. The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again, yielding:
$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
Remark 4.6 (Why SD =0=0 Is Unattainable).
The bounded-nonce constraint ($\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$) restricts the free parameter $\mathbf{t}$ to $\approx 2\gamma_{1}/|S|$ values per coefficient, whereas $\mathbb{Z}_{q}$-uniformity would require $q\approx 2^{23}$ values. The per-coefficient statistical distance from uniform is at least $1-2\gamma_{1}/(|S|\cdot q)>0$ for all practical parameters. Achieving SD =0=0 would require either unbounded nonces (incompatible with rejection sampling) or adding pairwise masks on $\mathbf{z}_{i}$ with $|S|\geq T+1$. The min-entropy guarantee is nonetheless strong: for $|S|\leq 17$, nonce entropy exceeds secret key entropy by $\geq 5\times$, ensuring robust share privacy without computational assumptions.
Corollary 4.7 (No Two-Honest Requirement in Coordinator-Based Profiles).
In Profiles P1 (TEE) and P3+ (trusted combiner), the protocol achieves response privacy with signing sets of size $|S|=T$ (minimum threshold). The two-honest assumption ($|S|\geq T+1$) is not needed: individual $\mathbf{W}_{i}$ values stay within the coordinator, and the min-entropy bound of Theorem 4.4 holds even with a single honest party. In Profile P2 (no coordinator), parties broadcast $\mathbf{W}_{i}$ directly; mask hiding (Lemma 4.9) protects these values under the condition $|S\setminus C|\geq 2$, which P2 already requires for r0-check share privacy.
Corollary 4.8 (Multi-Session Privacy).
Over $K$ signing sessions with fresh nonce DKGs, the total min-entropy of honest party $h$’s responses $(\mathbf{z}_{h}^{(1)},\ldots,\mathbf{z}_{h}^{(K)})$ is at least $K\cdot n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$ bits (each session contributes independently). For $|S|\leq 17$, this exceeds $K\cdot 24{,}000$ bits, providing robust multi-session privacy. We recommend periodic key rotation per standard operational practice.
Lemma 4.9 (Mask Hiding (for $\mathbf{V}_{i}$ and $\mathbf{W}_{i}$)).
@@
An honest party’s masked r0-check share $\mathbf{V}_{j}=\lambda_{j}c\mathbf{s}_{2,j}+\mathbf{m}_{j}^{(s2)}$ and masked commitment $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$ are computationally indistinguishable from uniform, given all information available to a coalition of <T<T parties, provided $|S\setminus C|\geq 2$.
Proof 4.10.
@@
Since $|S\setminus C|\geq 2$, party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{nonce})$ for some honest $k^{\prime}\neq j$. This seed is unknown to $\mathcal{A}$, so by PRF security, the term is indistinguishable from uniform. Thus $\mathbf{V}_{j}=(\text{fixed})+(\text{uniform-looking})\stackrel{{\scriptstyle c}}{{\approx}}\text{uniform}$. The same argument applies to $\mathbf{W}_{j}$.
Note: Mask hiding requires $|S\setminus C|\geq 2$. In P1 and P3+, individual $\mathbf{W}_{i}$ are sent only to the coordinator (Remark 3.7), so mask hiding for $\mathbf{W}_{i}$ is only needed in P2 where parties broadcast directly. The response $\mathbf{z}_{i}$ achieves statistical privacy with even $|S\setminus C|=1$ via the nonce DKG (Theorem 4.4), independent of mask hiding.
@@ 4.4 Nonce Distribution Security
Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})}+q_{S}\cdot\epsilon$ where the per-coordinate smooth Rényi divergence satisfies $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers. Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ with $n\ell=1280$ for ML-DSA-65. The tail probability satisfies $\epsilon<10^{-30}$.
The square root arises from the smooth Rényi divergence probability transfer (Cauchy-Schwarz): for any event $E$, $\Pr_{P}[E]\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot\Pr_{Q}[E]}+\epsilon$. The $q_{S}\cdot\epsilon$ term accounts for the smoothing cost across $q_{S}$ signing queries. Since the nonce coordinates are sampled independently, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the product of per-coordinate divergences. For ML-DSA-65 ($n\ell=1280$, $\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$, $\epsilon<2^{-100}$):
$\mathsf{Adv}_{\Pi}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot 2^{-192}}+q_{S}\cdot 2^{-100}$
$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}$ $R_{2}^{\mathsf{vec}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
$q_{S}=2^{30}$ $q_{S}=2^{64}$
$\leq 17$ 1.003 $\approx 2^{5.4}$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ $\approx 2^{-36}$
$\leq 25$ 1.014 $\approx 2^{25}$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ $\approx 2^{-36}$
$\leq 33$ 1.041 $\approx 2^{75}$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ $\approx 2^{-36}$
For $|S|\leq 17$, the $q_{S}\cdot\epsilon$ term dominates and the full-vector Rényi loss is only $\approx$3 bits. For $|S|\geq 25$, the $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant. This per-coordinate composition cost is inherent to the Rényi-based proof technique (shared with Raccoon [Raccoon2024]); tighter bounds via specialized Fiat-Shamir-with-Aborts reductions that exploit rejection sampling structure are an interesting open problem. Detailed calculations appear in Appendix D.
@@ 4.6 Concrete Security Bounds
Share Privacy ($\mathbf{z}_{i}$) Statistical (min-entropy $\geq 24$K bits) N/A
Mask Privacy ($\mathbf{V}_{i}$, $\mathbf{W}_{i}$) PRF 192 bits
Table 3: Concrete security of masked threshold ML-DSA-65. Parameters match NIST Security Level 3. Underlying M-SIS hardness is 192 bits; the proven threshold bound depends on $|S|$ via the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$: for $|S|\leq 17$ and $q_{S}=2^{30}$, $\mathsf{Adv}<2^{-70}$ (see Theorem 4.11). Share privacy is statistical (no computational assumptions); see Theorem 4.4.
The threshold construction inherits the M-SIS/M-LWE hardness of single-signer ML-DSA. The proven forgery bound includes a $\sqrt{\cdot}$ loss from Rényi divergence and a $q_{S}\cdot\epsilon$ term from Irwin-Hall nonces (Theorem 4.11); tightening these bounds via specialized reductions is an open problem shared with Raccoon [Raccoon2024].
@@ 4.7.1 Profile P1 UC Security
  • Round 2 Input: Receive masked commitments $\{(\mathbf{W}_{i},r_{i})\}_{i\in S}$ from parties (not broadcast)
[…]
  • Round 3 Input: Receive response shares $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$
@@
      Aggregate $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}\bmod q$ and $c\mathbf{s}_{2}=\sum_{i}\mathbf{V}_{i}\bmod q$
@@
  • Final Output: Send $(\tilde{c},\mathbf{z},\mathbf{h})$ or $\bot$ to all parties; never reveal $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$
@@
    $\mathcal{S}$ simulates honest parties’ Round 1 messages (commitments $\mathsf{Com}_{h}$) and Round 3 messages (responses $\mathbf{z}_{h}$, masked shares $\mathbf{V}_{h}$). Responses have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
@@
[…]
@@
    $\mathcal{F}_{\mathsf{TEE}}$ computes the r0-check and hint internally, outputs only the final signature (or $\bot$).
Since $\mathcal{F}_{\mathsf{TEE}}$ never reveals $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$, the responses $\mathbf{z}_{h}$ have high min-entropy (Theorem 4.4, statistical), and the adversary’s view contains only $(\mathbf{w}_{1},\tilde{c},\sigma)$—all implicit in the signature—the real and ideal executions are indistinguishable. The distinguishing advantage is bounded by $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (for the mask-based values only; the response privacy is statistical and contributes negligible statistical distance).
@@ 4.7.2 Profile P3+ Semi-Async Security Model
Theorem 4.17 (Profile P3+ UC Security (Semi-Async)).
@@ 4.7.3 Summary of UC Security
  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.13). EUF-CMA follows from Theorem 4.2; statistical share privacy from Theorem 4.4. Achieves 3 online rounds plus 1 offline preprocessing round (nonce DKG).
@@ 5.2 Benchmark Methodology
Our Rust implementation ($\approx$5500 lines of core protocol logic, plus $\approx$4000 lines of tests and utilities) provides optimized performance through parallel polynomial operations using Rayon, NTT-based polynomial multiplication with Montgomery reduction, and efficient EdaBit and Beaver triple pooling with lazy generation.
@@
All measurements are per-attempt: a single execution of the online signing protocol (3 rounds for P1, 5 for P2, 2 for P3+), which either succeeds (outputs $\sigma$) or aborts (due to rejection sampling). Nonce DKG preprocessing is excluded from per-attempt timing, as it is performed offline.